# Magnetism in Twisted Bilayer WSe2

### v1

@@ Abstract
Using a self-consistent Hartree-Fock theory, we show that the recently observed ferromagnetism in twisted bilayer WSe2 [Nat. Commun. 16, 1959 (2025)] can be understood as a Stoner-like instability of interaction-renormalized moiré bands. We quantitatively reproduce the observed Lifshitz transition as function of hole filling and applied electric field that marks the boundary between layer-hybridized and layer-polarized regimes. The former supports a ferromagnetic valley-polarized ground state below half-filling, developing a topological charge gap at half-filling for small twists. At larger twist angles there is a transition to a gapped triangular Néel antiferromagnet. The layer-polarized regime supports a stripe antiferromagnet below half-filling and a wing-shaped multiferroic ground state above half-filling. We map the evolution of these states as a function of filling factor, electric field, twist angle, and interaction strength. Beyond providing an understanding of recent experiments, our methodology is applicable to a broad class of moiré systems.
Moiré superlattice engineering of small rotational mismatches between atomically thin layers has opened up new avenues for studying strongly-correlated electronic phases in two-dimensional systems. In graphene-based moirés, such as magic-angle twisted bilayer graphene, electron-electron interactions in the nearly flat moiré bands near charge neutrality give rise to a rich phenomenology including correlated insulators [1], superconductivity [2], among others. More recently, there has been a lot of experimental progress on twisted transition metal dichalcogenides (TMDs), where strong spin-orbit coupling, topology, and in-situ moiré band engineering provide new pathways for realizing correlated quantum phases [3, 4, 5]. Here we focus on twisted bilayer WSe2 (tWSe2) where recent experiments have reported correlated magnetic states [6] and superconductivity [7, 8].
@@
In this Letter, we address the first two experimental observations by incorporating long-range Coulomb interactions within a Hartree-Fock (HF) scheme. The third feature, namely the observation of superconductivity at both small [7] and large [8] electric fields, has been discussed in some very recent theoretical works [10, 11, 12, 13, 14, 15, 16, 17]. Our main results are summarized in Fig. 1. We demonstrate that both the experimentally observed phase boundary and triangular ferromagnetic region can be understood within a self-consistent mean-field theory. We interpret our findings in terms of a Stoner instability of interaction-renormalized moiré bands of the symmetry-unbroken HF state. The success of our approach suggests a more broadly applicable hierarchy of approximations, where the theory of symmetry-broken phases should be based on an unbroken parent state with strongly renormalized moiré bands. To guide future experiments, we further investigate the phase diagram as a function of hole filling and electric field, as well as the nature of the HF ground state at half-filling.
Model — Our starting point is the Hamiltonian $H=H_{0}+H_{\text{int}}$ where $H_{0}$ is the noninteracting low-energy moiré theory for the valence band near valley $K/K^{\prime}$. Due to strong Ising spin-orbit coupling, spin $S_{z}$ is locked to the valley and $H_{0}=\sum_{\sigma=\uparrow,\downarrow}\int d^{2}\bm{r}\,\psi_{\sigma}^{\dagger}(\bm{r})\mathcal{H}_{0}^{\sigma}\psi_{\sigma}(\bm{r})$ with
@@
in the layer basis within the local-stacking approximation (LSA) [18, 19]. Here we take $m_{*}=0.45m_{e}$ for the effective mass, $V_{z}$ is an interlayer bias from the applied electric field, and $\varepsilon(\bm{r})$ [$t(\bm{r})$] are intralayer [interlayer] moiré potentials. These can be written as
@@
where the first sum runs over monolayer reciprocal vectors $\bm{b}$. Here $\bm{b}_{j}^{(2,3)}=R(\pm 2\pi/3)\bm{b}_{j}^{(1)}$ and $\bm{K}=4\pi\hat{x}/3a$ with $a=3.317$. For example, for the first star: $\bm{b}_{1}^{(1)}=4\pi\hat{y}/\sqrt{3}a$ and $|\bm{K}+\bm{b}|=4\pi/3a$. Here $\bm{\phi}(\bm{r})=(a/a_{M})\hat{z}\times\bm{r}+\bm{u}(\bm{r})$ is the local stacking with moiré length $a_{M}=a/[2\sin(\theta/2)]$ and $\bm{u}(\bm{r})$ is the displacement field due to lattice relaxation [20, 21]. The latter effectively modifies the rigid lattice moiré potentials in Eqs. (2) and (3). We find that lattice relaxation becomes significant for twist angles $\theta<3^{\circ}$. In this work, we consider up to three stars of reciprocal lattice vectors in the expansion in Eqs. (2) and (3) using parameters obtained from density-functional theory (DFT) for untwisted bilayers. Computational details are in the Supplemental Material (SM)  [^1] See Supplemental Material [url] for … and nonzero values are shown in Table 1.
@@
  $(6.61,89^{\circ})$ $(0.21,-94^{\circ})$ $12.96$ $-1.95$ $-0.58$
$\sqrt{3}$ $2\sqrt{3}$ $1$ $2$ $\sqrt{7}$
@@
For the twist angles we consider, the long-wavelength moiré leaves valleys decoupled, and the system has U(1) valley charge conservation, i.e., $[H,S_{z}]=0$. In a single valley, the remaining symmetries are the magnetic point group $D_{3}(C_{3})=\left<\mathcal{C}_{3z},\mathcal{C}_{2y}\mathcal{T}\right>$ where $\mathcal{T}=i\sigma_{y}K$ is time reversal, and moiré translations. In addition, for $V_{z}=0$ the moiré bands are spin degenerate: $E_{n}^{\sigma}(\bm{k})\overset{\mathcal{P}}{=}E_{n}^{\sigma}(-\bm{k})\overset{\mathcal{T}}{=}E_{n}^{-\sigma}(\bm{k})$. Here $\mathcal{P}:(\bm{r}\mapsto-\bm{r},\tau_{x})$ where $\tau_{x}$ flips the layers, is an “intravalley inversion” symmetry built into Eq. (1) by the LSA. It is only exact in the first star of the moiré potentials and is broken when either $V_{z}\neq 0$ (which also breaks $\mathcal{C}_{2y}$) or we go beyond LSA  [^2] In this case $t^{*}(\bm{r})\neq t(-\bm{r})$ and we only have $\varepsilon_{2}(x,y)=\varepsilon_{1}(-x,y)$ by $\mathcal{C}_{2y}\mathcal{T}$ symmetry. For example, $\mathcal{P}$ symmetry is broken by $\arg(w_{1}/w_{2})\neq 0,\pi$ and $\psi_{2}\neq 0,\pi$ which is the phase of the second star that is equal for both layers.. The latter lifts the spin degeneracy of the moiré bands, which can be seen along the $\gamma-m$ line in DFT calculations [24, 25].
@@
Hartree-Fock parent state — We first consider the symmetry-unbroken HF state. While this is not necessarily the ground state, it serves as a parent state for symmetry-broken phases. In contrast to twisted bilayer graphene [27, 28] and graphene on h-BN [29] where away from charge neutrality band-renormalization is dominated by Hartree term, we find that for tWSe2 the self-energy is always dominated by Fock term. Hence the same exchange interactions responsible for magnetic order also dominate band renormalization. In Fig. 1 we show the density of states (DOS) of the single-particle moiré bands (a) and the parent unbroken HF state (b) for $\theta=3.65^{\circ}$ as function of hole doping $\nu$ and interlayer bias $V_{z}$. We find that the DOS is significantly modified by Coulomb interactions. This reshapes the VHS and softens the Lifshitz transition between layer-hybridized and layer-polarized regimes [Fig. 1(c)] as experimentally observed [6]. In Fig. 1(d) we compare the Lifshitz transition of the parent state to the experimental data of Ref. [6]. We find that our results can reproduce the data for $\epsilon\approx 30$ where we assume that the interlayer bias is linear in the applied electric field. From the experimental geometry, one would expect $\epsilon\sim 4$; however, on very general grounds, we can anticipate a strong suppression of the Fock interactions in moiré materials that is captured phenomenologically using a larger effective dielectric constant [30]. We note that other theory works on TMDs also use larger values for the dielectric constant [31, 32, 33].
@@
Figure 2: Ground state magnetic order and layer polarization for $\theta=3.65^{\circ}$ and $\epsilon=25$. (a) Out-of-plane spin polarization corresponding to the FMz phase, which exists both near charge neutrality and in regions where Fock interactions enhance the DOS [Fig. 1(b)]. The experimentally observed ferromagnetic region at small $V_{z}$ below half-filling ($\nu=1$) is clearly visible. (b) In-plane spin polarization of the IVC stripe antiferromagnet, which exists only in a small $\nu$ window in layer-polarized regions. (c) Order parameter of the $120^{\circ}$ AFM: $\sum_{\bm{k}}f_{\bm{k}}\langle c_{\bm{k},\uparrow}^{\dagger}c_{\bm{k}+\bm{q},\downarrow}\rangle$. (d) Layer polarization where dashed lines mark the FMz wing-shaped region from (a).
[…]
@@
Magnetic orders — Next, we consider three symmetry-breaking magnetic orders: (i) a spin and valley polarized ferromagnetic state (FMz) with finite $S_{z}$ polarization; (ii) an intervalley coherent (IVC) antiferromagnetic state that conserves moiré translations (AFMxy), and (iii) an IVC triangular Néel AFM that has a $\sqrt{3}\times\sqrt{3}$ reconstruction of the moiré cell (120 AFM). The AFMxy state is an uncompensated stripe antiferromaget while the 120 AFM hosts clockwise rotating spins for $V_{z}>0$, and anticlockwise for $V_{z}<0$. While all the IVC orders break $\mathcal{T}$ and valley U(1), they conserve an effective time-reversal symmetry $\mathcal{T}^{\prime}=e^{i\pi\sigma_{z}/2}\mathcal{T}$ with $(\mathcal{T}^{\prime})^{2}=1$ [35, 36].
@@
In Fig. 2 we show the three magnetic order parameters in the ground state, for $\theta=3.65^{\circ}$ as a function of filling and interlayer bias. These are $S_{z}$, $S_{x,y}$, and $\langle c_{\bm{k},\uparrow}^{\dagger}c_{\bm{k}+\bm{q},\downarrow}\rangle$ with $\bm{q}=\pm 4\pi\hat{y}/(3L)$. We find the FMz phase [Fig. 2(a)] exhibits a fully-polarized phase near $V_{z}=0$ below half-filling, and two “wings” with partial polarization at finite $V_{z}$. This can be understood in terms of Stoner-like ferromagnetism [37] of the HF parent state, where band renormalization from Fock interactions warps the VHS and induces a triangular region with enhanced DOS below half-filling near zero interlayer bias. These features are consistent with recent MCD measurements on tWSe2 [6] and similar observations in tMoTe2 [3]. We also find FMz at small filling, which may be invisible in MCD due to the low carrier density. On the other hand, the uncompensated stripe AFMxy has a small in-plane magnetization [Fig. 2(b)] and appears at finite interlayer bias below half-filling, while the $120^{\circ}$ AFM state [Fig. 2(c)] emerges at near half-filling in a finite $V_{z}$ window. Moreover, we find this state has a finite charge gap at $\nu=1$ as observed in experiment [7, 6]. We further show the ground-state layer polarization in Fig. 2(d). We see that below half filling the FMz state is layer hybridized while the wings have a finite layer polarization giving rise to a multiferroic state. While the stripe AFMxy is mostly layer polarized, the $120^{\circ}$ AFM phase shows a reduced layer polarization compared to the unbroken state.
The complete $\nu$$V_{z}$ phase diagram for $\theta=3.65^{\circ}$ and $\epsilon=25$ is shown in Fig. 4(a). Near charge neutrality, the system behaves as a 2DEG for which the Stoner mechanism gives full $S_{z}$ polarization [22]. As we add holes, the FMz phase emerges in regions of enhanced DOS of the unbroken state in accordance to the Stoner mechanism, while it gives way to a $\sqrt{3}\times\sqrt{3}$ spin density instability near half filling anticipated from the spin susceptibility [13] giving rise to the $120^{\circ}$ AFM state. On the other hand, the IVC stripe AFMxy is favored compared to the FMz phase by the spin splitting of the dispersion which is driven by the applied field $V_{z}$ [14, 38] and hence it only appears at finite interlayer bias.
Figure 3: Evolution of DC Hall conductivity $\sigma_{xy}$, which characterizes the degree of $\mathcal{T}$ breaking, as a function of interlayer bias $V_{z}$ and hole filling factor $\nu$ for different values of the dielectric constant $\epsilon$. The “wing” feature emerges as a consequence of strong correlations. Contrary to $S_{z}\right>$, we find that $\sigma_{xy}$ is a good probe for the layer-polarized FMz phase.
@@
Figure 4: Phase diagram and ground state at half filling versus twist angle and interaction strength. (a) Phase diagram for $\theta=3.65^{\circ}$ and $\epsilon=25$. The $\nu=1$ phase diagram versus (b) interaction strength ($1/\epsilon$) and (c) twist angle, where the color gives the charge gap $\Delta$. The gapped IVC $120^{\circ}$ AFM persist even for weak interactions in regions where the Fermi energy is close to the VHS on the unbroken parent state (M).
DC Hall conductivity — The Hall conductivity $\sigma_{xy}$ serves as an important experimental observable in transport that can probe the $\mathcal{T}$ breaking of the FMz phase [7]. We anticipate that $\sigma_{xy}$ will soon be measured in this system both as a function of filling and interlayer bias. The Hall conductivity can be expressed as
$\sigma_{xy}=\frac{e^{2}}{\hbar}\sum_{n}\int_{\text{MBZ}}\frac{d^{2}\bm{q}}{(2\pi)^{2}}f_{n,\bm{q}}\Omega_{n,\bm{q}},$ (5)
where the sum runs over HF bands, $f_{n,\bm{q}}$ is the occupation and $\Omega_{n,\bm{q}}$ is the Berry curvature. In Fig. 3, we show $\sigma_{xy}$ in the $\nu$$V_{z}$ plane for different $\epsilon$. The symmetry-unbroken and IVC states have vanishing $\sigma_{xy}$ due to $\mathcal{T}^{\prime}$. As a result, $\sigma_{xy}$ only probes the FMz phase. Note that while we find large $\sigma_{z}\right>$ in Fig. 2(b) near charge neutrality, the Hall conductivity is absent because both the density of holes and Berry curvature is small. Moreover, for $V_{z}=0$, regardless of the interaction strength, we find finite $\sigma_{xy}$ only for $\nu<1$, consistent with experiment [6]. Note that $\sigma_{xy}$ is absent at half-filling because the gap is trivial and the $120^{\circ}$ AFM has $\mathcal{T}^{\prime}$ symmetry. Contrary to the out-of-plane magnetization, we find that $\sigma_{xy}$ is a good probe for the layer-polarized FMz.
@@
Correlated insulator at half-filling — At half-filling, transitions between a superconductor, a correlated insulator, and a metal were recently reported [7]. To understand the origin of these phases in WSe2, we examine the ground state in the self-consistent HF approximation. Our results at half-filing are shown here for completeness and are in agreement with previous reports in the literature focusing mostly on tMoTe2 [31, 32, 39, 33, 40]. We emphasize that the filling dependence of these states has not been considered previously. In Fig. 4(b) and (c), we show the phase diagram and charge gap as a function of the interaction strength, interlayer bias, and twist angle. For small interlayer bias $V_{z}$, as we increase the interaction strength, the metallic symmetry-unbroken phase (M) gives way to a $120^{\circ}$ AFM with a topologically trivial gap. The experimentally observed superconductivity [7, 8] is believed to be mediated by magnons of the $120^{\circ}$ AFM. At larger $V_{z}$ the single-particle bandwidth increases yielding a metallic phase except for strong interactions. However, for intermediate $V_{z}$, the $120^{\circ}$ AFM persists even for weak interactions. This is due to the proximity of the Fermi energy to the VHS of the unbroken state. The FMz phase only becomes the ground state for very strong interactions yielding a quantum anomalous Hall insulator with unit Chern number. For fixed $\epsilon=25$ the FMz state appears for small twists close to $V_{z}=0$ transitioning to a trivially gapped stripe AFM as $V_{z}$ increases. For larger twists, a gapped $120^{\circ}$ AFM dominates for small interlayer bias until $|V_{z}|=25~{}\text{meV}$ where the gap closes to a the unbroken phase due to the increased single-particle bandwidth.

### v2

# Itinerant Magnetism in Twisted Bilayer WSe2 and MoTe2

@@ Abstract
Using a self-consistent Hartree-Fock theory, we show that the recently observed ferromagnetism in twisted bilayer WSe2 [Nat. Commun. 16, 1959 (2025)] can be understood as a Stoner-like instability of interaction-renormalized moiré bands. We quantitatively reproduce the observed Lifshitz transition as function of hole filling and applied electric field that marks the boundary between layer-hybridized and layer-polarized regimes. The former supports a ferromagnetic valley-polarized ground state below half-filling, developing a topological charge gap at half-filling for smaller twist angles. At larger twist angles, the system hosts a gapped triangular Néel antiferromagnet. On the other hand, the layer-polarized regime supports a stripe antiferromagnet below half-filling and a wing-shaped multiferroic ground state above half-filling. We map the evolution of these states as a function of filling factor, electric field, twist angle, and interaction strength. Our results demonstrate that long-range exchange in a symmetry-unbroken parent state with strongly renormalized moiré bands provides a broadly applicable framework to understand itinerant magnetism in moiré TMDs.
Moiré superlattice engineering of small rotational mismatches between atomically thin layers has opened up new avenues for studying strongly-correlated electronic phases in two-dimensional systems. In graphene-based moirés, such as magic-angle twisted bilayer graphene (TBG), electron-electron interactions in the nearly flat moiré bands near charge neutrality give rise to a rich phenomenology including correlated insulators [1], superconductivity [2], among others. More recently, there has been a lot of experimental progress on twisted transition metal dichalcogenides (TMDs), where strong spin-orbit coupling, topology, and in-situ moiré band engineering provide new pathways for realizing correlated quantum phases [3, 4, 5]. Here we focus on twisted bilayer WSe2 (tWSe2) where recent experiments have reported correlated magnetic states [6, 7] and superconductivity [8, 9].
@@
In this Letter, we address the first two experimental observations by incorporating long-range Coulomb interactions within a Hartree-Fock (HF) scheme. The third feature, namely the observation of superconductivity at both small [8] and large [9] electric fields, has been discussed in some very recent theoretical works [11, 12, 13, 14, 15, 16, 17, 18, 19]. Our main results are summarized in Fig. 1. We demonstrate that both the experimentally observed phase boundary and triangular ferromagnetic region can be understood within a self-consistent mean-field theory. We interpret our findings in terms of a Stoner instability of interaction-renormalized moiré bands of the symmetry-unbroken HF state. The success of our approach suggests a more broadly applicable hierarchy of approximations, where the theory of symmetry-broken phases should be based on an unbroken parent state with strongly renormalized moiré bands. To guide future experiments, we further investigate the phase diagram and orbital magnetization as a function of hole filling and electric field, interaction-driven Lifshitz transitions, as well as the nature of the HF ground state at half-filling.
Model — Our starting point is the Hamiltonian $\mathcal{H}=\mathcal{H}_{0}+\mathcal{H}_{\text{int}}$ where $\mathcal{H}_{0}$ is the noninteracting low-energy moiré theory for the valence band near valley $K/K^{\prime}$. Due to strong Ising spin-orbit coupling, spin $S_{z}$ is locked to the valley and $\mathcal{H}_{0}=\sum_{\sigma=\uparrow,\downarrow}\int d^{2}\bm{r}\,\psi_{\sigma}^{\dagger}(\bm{r})\mathcal{H}_{0}^{\sigma}\psi_{\sigma}(\bm{r})$ with
@@
in the layer basis within the local-stacking approximation (LSA) [20, 21]. Here we take $m_{*}=0.45m_{e}$ for the effective mass, $V_{z}$ is an interlayer bias from the applied electric field, and $\varepsilon(\bm{r})$ $t(\bm{r})$] are intralayer [interlayer] moiré potentials. These can be written as
@@
where the first sum runs over monolayer reciprocal vectors $\bm{b}$. Here $\bm{b}_{j}^{(2,3)}=R(\pm 2\pi/3)\bm{b}_{j}^{(1)}$ and $\bm{K}=4\pi\hat{x}/3a$ with $a=3.317$ Å. For example, for the first star: $\bm{b}_{1}^{(1)}=4\pi\hat{y}/\sqrt{3}a$ and $|\bm{K}+\bm{b}|=4\pi/3a$. Here $\bm{\phi}(\bm{r})=(a/a_{M})\hat{z}\times\bm{r}+\bm{u}(\bm{r})$ is the local stacking with moiré length $a_{M}=a/[2\sin(\theta/2)]$ and $\bm{u}(\bm{r})$ is the displacement field due to lattice relaxation [22, 23]. The latter effectively modifies the rigid lattice moiré potentials in Eqs. (2) and (3). We find that lattice relaxation becomes significant for twist angles $\theta<3^{\circ}$. In this work, we consider up to three stars of reciprocal lattice vectors in the expansion in Eqs. (2) and (3) using parameters obtained from density-functional theory (DFT) for untwisted bilayers. Computational details are in the Supplemental Material (SM)  [^1] See Supplemental Material [url] for … and nonzero values are shown in Table 1.
@@
  $(6.61,89^{\circ})$ $(0.21,-94^{\circ})$ $12.96$ 1.95-1.95 0.58-0.58
$\sqrt{3}$ $2\sqrt{3}$ 11 22 $\sqrt{7}$
@@
For the twist angles we consider, the long-wavelength moiré leaves valleys decoupled, and the system has U(1) valley charge conservation, i.e., $\mathcal{H},S_{z}]=0$. In a single valley, the remaining symmetries are the magnetic point group $D_{3}(C_{3})=\left<\mathcal{C}_{3z},\mathcal{C}_{2y}\mathcal{T}\right$> where $\mathcal{T}=i\sigma_{y}K$ is time reversal, and moiré translations. In addition, for $V_{z}=0$ the moiré bands are spin degenerate: $E_{n}^{\sigma}(\bm{k})\overset{\mathcal{P}}{=}E_{n}^{\sigma}(-\bm{k})\overset{\mathcal{T}}{=}E_{n}^{-\sigma}(\bm{k})$. Here $\mathcal{P}:(\bm{r}\mapsto-\bm{r},\tau_{x})$ where $\tau_{x}$ flips the layers, is an “intravalley inversion” symmetry built into Eq. (1) by the LSA. It is only exact in the first star of the moiré potentials and is broken when either $V_{z}\neq 0$ (which also breaks $\mathcal{C}_{2y}$) or we go beyond LSA  [^2] In this case $t^{*}(\bm{r})\neq t(-\bm{r})$ and we only have $\varepsilon_{2}(x,y)=\varepsilon_{1}(-x,y)$ by $\mathcal{C}_{2y}\mathcal{T}$ symmetry. For example, $\mathcal{P}$ symmetry is broken by $\arg(w_{1}/w_{2})\neq 0,\pi$ and $\psi_{2}\neq 0,\pi$ which is the phase of the second star that is equal for both layers.. The latter lifts the spin degeneracy of the moiré bands, which can be seen along the $\gamma-m$ line in DFT calculations [26, 27].
@@
Hartree-Fock parent state — We first consider the symmetry-unbroken HF state. While this is not necessarily the ground state, it serves as a parent state for symmetry-broken phases. In contrast to twisted bilayer graphene [29, 30] and graphene on h-BN [31] where away from charge neutrality band-renormalization is dominated by Hartree term, we find that for tWSe2 the self-energy is always dominated by Fock term. Hence the same exchange interactions responsible for magnetic order also dominate band renormalization. In Fig. 1 we show the density of states (DOS) of the single-particle moiré bands (a) and the parent unbroken HF state (b) for $\theta=3.65^{\circ}$ as function of hole doping $\nu$ and interlayer bias $V_{z}$. We find that the DOS is significantly modified by Coulomb interactions. This reshapes the VHS and softens the Lifshitz transition between layer-hybridized and layer-polarized regimes [Fig. 1(c)] as experimentally observed [7]. In Fig. 1(d) we compare the Lifshitz transition of the parent state to the experimental data of Ref. [7]. We find that our results can reproduce the data for $\epsilon\approx 30$ where we assume that the interlayer bias is linear in the applied electric field. From the experimental geometry, one would expect $\epsilon\sim 4$; however, on very general grounds, we can anticipate a strong suppression of the Fock interactions in moiré materials that is captured phenomenologically using a larger effective dielectric constant. We note that other theory works on TMDs also use larger values for the dielectric constant [32, 33, 34].
@@
Figure 2: Ground state magnetic order and layer polarization for $\theta=3.65^{\circ}$ and $\epsilon=25$. (a) Out-of-plane spin polarization corresponding to the FMz phase, which exists both near charge neutrality and in regions where Fock interactions enhance the DOS [Fig. 1(b)]. The experimentally observed ferromagnetic region at small $V_{z}$ below half-filling ($\nu=1$) is clearly visible. (b) In-plane spin polarization of the IVC stripe antiferromagnet, which exists only in a small $\nu$ window in layer-polarized regions. (c) Order parameter of the $120^{\circ}$ AFM: $\sum_{\bm{k}}f_{\bm{k}}\langle c_{\bm{k},\uparrow}^{\dagger}c_{\bm{k}+\bm{q},\downarrow}\rangle$. (d) Order parameter of the generalized Wigner crystal state with $\sqrt{3}\times\sqrt{3}$, $2\times 2$, and $2\times 3$ reconstructions of the moiré unit cell.
Some features of our numerical results can be understood from a toy model that neglects the moiré potentials. In this scenario, we have two parabolic bands $E^{\mathrm{NI}}_{\bm{k}}$ and $E^{\mathrm{NI}}_{\bm{k}}-V_{z}$ with $E^{\mathrm{NI}}_{\bm{k}}=-k^{2}/2m_{*}$ and $V_{z}>0$, as illustrated in Fig. 1(e). A Lifshitz transition occurs when the Fermi energy reaches the top of the second band, i.e, for $\nu=2\Omega_{M}\int_{-V_{z,c}}^{0}\mathrm{DOS}(E)\,dE$, with $\Omega_{M}$ the moiré cell area, and the factor 2 accounts for spin. Consequently, we obtain $V_{z,c}=\pi\nu/\Omega_{M}m^{*}$ which yields the linear boundary in the single-particle limit as shown in Fig. 1(a). Adding Coulomb interactions renormalizes the top band: $E^{\text{HF}}_{\bm{k}}=E^{\text{NI}}_{\bm{k}}+\Sigma^{\text{F}}_{\bm{k}}$, where $\Sigma^{\text{F}}_{\bm{k}}=\int d^{2}\bm{q}V_{\bm{k}+\bm{q}}f_{\bm{q}}/(2\pi)^{2}$  [^3] For a uniform electron gas the Hartree term is canceled by the jellium background is the Fock self-energy with $f_{\bm{q}}$ the occupation. The integral can be carried out analytically [24] and $E^{\text{HF}}_{\bm{k}}\simeq(e^{2}k_{\text{F}}/\epsilon)\left[1-(k/2k_{\text{F}})^{2}\right]$ where $k_{\text{F}}=\sqrt{2\pi\nu/\Omega_{M}}$. Hence the Fock correction shifts the upper band upward and reduces the mass $1/m_{*}^{\text{F}}=1/m_{*}+e^{2}/2\epsilon k_{\mathrm{F}}$. For a given filling, the critical bias is now determined by
@@
Magnetic orders — Next, we consider three symmetry-breaking magnetic orders: (i) a spin and valley polarized ferromagnetic state (FMz) with finite $S_{z}$ polarization; (ii) an intervalley coherent (IVC) antiferromagnetic state that conserves moiré translations (AFMxy); (iii) an IVC triangular Néel AFM that has a $\sqrt{3}\times\sqrt{3}$ reconstruction of the moiré cell (120 AFM), and (iv) a generalized Wigner crystal (GWC) with $\sqrt{3}\times\sqrt{3}$, $2\times 2$, and $2\times 3$ reconstructions of the moiré unit cell, which realizes a valley-polarized spin-density wave. The AFMxy state is an uncompensated stripe antiferromaget while the 120 AFM hosts clockwise rotating spins for $V_{z}>0$, and anticlockwise for $V_{z}<0$. While all the IVC orders break $\mathcal{T}$ and valley U(1), they conserve an effective time-reversal symmetry $\mathcal{T}^{\prime}=e^{i\pi\sigma_{z}/2}\mathcal{T}$ with $(\mathcal{T}^{\prime})^{2}=1$ [36, 37].
@@
In Fig. 2 we show the three magnetic order parameters in the ground state, for $\theta=3.65^{\circ}$ as a function of filling and interlayer bias. These are $S_{z}$, $S_{x,y}$, and $\langle c_{\bm{k},\uparrow}^{\dagger}c_{\bm{k}+\bm{q},\downarrow}\rangle$ with $\bm{q}=\pm 4\pi\hat{y}/(3L)$. We find the FMz phase [Fig. 2(a)] exhibits a fully-polarized phase near $V_{z}=0$ below half-filling, and two “wings” with partial polarization at finite $V_{z}$. This can be understood in terms of Stoner-like ferromagnetism [38] of the HF parent state, where band renormalization from Fock interactions warps the VHS and induces a triangular region with enhanced DOS below half-filling near zero interlayer bias. These features are consistent with recent MCD measurements on tWSe2 [7] and similar observations in tMoTe2 [3]. We also find FMz at small filling, which may be invisible in MCD due to the low carrier density. On the other hand, the uncompensated stripe AFMxy has a small in-plane magnetization [Fig. 2(b)] and appears at finite interlayer bias below half-filling, while the $120^{\circ}$ AFM state [Fig. 2(c)] emerges at near half-filling in a finite $V_{z}$ window. Moreover, we find this state has a finite charge gap at $\nu=1$ as observed in experiment [8, 7]. In addition, we find generalized Wigner crystal states as the ground states near $1/3$, $1/4$, and $1/6$ filling (one hole per reconstructed cell), and they remain stable under finite interlayer bias. This trend is consistent with experimental observations near $3.5^{\circ}$ twist [7]. Allowing translational symmetry breaking also produces insulating states at $\nu=1/2$, $2/3$, and $3/4$, although these are much weaker with only a slight energy gain relative to the competing unreconstructed ferromagnet [24]. For completeness, we also compute the phase diagram for tMoTe2 at $\theta=3.5^{\circ}$ [24], where the main difference is that the FMz state becomes the ground state at half filling matching recent experiments [39].
Figure 3: Orbital magnetization per moiré unit cell. Results are shown as a function of filling $\nu$ and interlayer potential $V_{z}$ for the FMz phase for tWSe2 at $\theta=3.65^{\circ}$ (left) and tMoTe2 at $\theta=3.5^{\circ}$ (right). Twisted WSe2 peaks at about $0.4~\mu_{B}$ near half filling, while twisted MoTe2 shows a stronger response of about $1.2~\mu_{B}$ that is further enhanced by finite $V_{z}$, reflecting the larger Berry curvature and narrower bands.
@@
Figure 4: Phase diagram, Fermi surface topology and ground state at half filling versus twist angle and interaction strength. (a) Phase diagram for $\theta=3.65^{\circ}$ and $\epsilon=25$, shown together with a representative Fermi surface of the gapless phases. Red and blue indicate the Fermi surfaces of the two valleys. The $\nu=1$ phase diagram versus (b) interaction strength ($1/\epsilon$) and (c) twist angle, where the color gives the charge gap $\Delta$. The gapped IVC $120^{\circ}$ AFM persist even for weak interactions in regions where the Fermi energy is close to the VHS on the unbroken parent state (M).
[…]
Spin and orbital magnetization — To investigate the origin of the observed magnetization, we compare spin and orbital contributions. The spin contribution is given by the out-of-plane polarization shown in Fig. 2(a), and reaches about $1\ \mu_{B}$ per moiré cell near half filling. The orbital part [24, 40, 41] is shown in Fig. 3(a). We find that tMoTe2 has a much larger orbital response than tWSe2, with a peak of about $1.2\ \mu_{B}$ per moiré cell slightly below half filling, which is further enhanced for finite $V_{z}$. This reflects larger Berry curvature and narrower valley Chern bands. Unlike TBG at odd fillings—where spin polarization is weak [42], the magnetism in tWSe2 and tMoTe2 has comparable spin and orbital parts due to spin–valley locking. The moiré-band contributions match DFT-based estimates [43] but are a factor of $\sim$2 lower than experiments [44]. As discussed in Refs. [45, 46], this discrepancy may stem from the fact that our continuum model is constructed from $d_{x^{2}-y^{2}}\pm id_{xy}$ orbitals rather than the full Bloch wave functions.
@@
Correlated insulator at half-filling — At half-filling, transitions between a superconductor, a correlated insulator, and a metal were recently reported [8]. To understand the origin of these phases in WSe2, we examine the ground state in the self-consistent HF approximation. Our results at half-filing are shown here for completeness and are in agreement with previous reports in the literature focusing mostly on tMoTe2 [32, 33, 47, 34, 48]. We emphasize that the filling dependence of these states has not been considered previously. In Fig. 4(b) and (c), we show the phase diagram and charge gap as a function of the interaction strength, interlayer bias, and twist angle. For a small interlayer bias $V_{z}$, as we increase the interaction strength, the metallic symmetry-unbroken phase (M) gives way to a $120^{\circ}$ AFM with a topologically trivial gap. The experimentally observed superconductivity [8, 9] is believed to be mediated by magnons of the $120^{\circ}$ AFM. At larger $V_{z}$ the single-particle bandwidth increases yielding a metallic phase except for strong interactions. However, for intermediate $V_{z}$, the $120^{\circ}$ AFM persists even for weak interactions. This is due to the proximity of the Fermi energy to the VHS of the unbroken state. The FMz phase only becomes the ground state for very strong interactions yielding a quantum anomalous Hall insulator with unit Chern number. For fixed $\epsilon=25$ the FMz state appears for small twists close to $V_{z}=0$ meV transitioning to a trivially gapped stripe AFM as $V_{z}$ increases. For larger twists, a gapped $120^{\circ}$ AFM dominates for small interlayer bias until $|V_{z}|=25~\text{meV}$ where the gap closes to a the unbroken phase due to the increased single-particle bandwidth. For small twist angles the FMz and AFMxy states differ in energy by less than $0.1$ meV per moiré cell [24], making the ground state sensitive to strain and other small perturbations. In this regime the FM–AFM balance at half filling is less reliable, and a weak out-of-plane magnetic field ($B\gtrsim 0.3$ T) is needed to stabilize the FMz state in experiments.